OpenStudy (anonymous):

Part 1: Create a scenario for an arithmetic sequence. For example, Jasmine practices the piano for ______ minutes on Monday. Every day she ___________ her practice time by _________. If she continues this pattern, how many minutes will she practice on the 7th day? Be sure to fill in the blanks with the words that will create an arithmetic sequence. Use your scenario to write the function for the 7th term in your sequence using sequence notation. Part 2: Create a scenario for a geometric sequence. For example, Anthony goes to the gym for ______ minutes on Monday. Every day he _______his gym

OpenStudy (anonymous):

Part 3: Use your scenario from part 2 to write a question that will lead to using the geometric series formula. Use the formula to solve for Sn in your scenario.

OpenStudy (anonymous):

I just need help with part 3... I got my formula I just dont understand how to write a question.

OpenStudy (anonymous):

@solomon_zelman

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@team_cuban

OpenStudy (anonymous):

@team_cuban

OpenStudy (anonymous):

What did you put so far?

OpenStudy (anonymous):

Anthony goes to the gym for 10 minutes on Monday. Every day he increases his gym time by doubling the minutes of his last work out. He will spend 160 minutes at the gym on the 5th day. a^n=a^1*r^n-1 7=10*2^7-1

OpenStudy (anonymous):

s^7=10-10(2^7) / 1-7

OpenStudy (anonymous):

Okay now write a question using the formula.

OpenStudy (anonymous):

I have no idea how to do that. The question was already written, that is why they asked me to make a formula. Sorry I am bad at this.

OpenStudy (anonymous):

?? @team_cuban